OpenStudy (anonymous):

During World War 2, there were some cases where the crew fell out of burning aircraft without a parachute and survived the fall. Assume that the crew member reached a constant terminal speed of 127.6 km/hr prior to hitting a stack of loose hay. If the crew member can survive an acceleration of 35.0 g, where g is the gravitational constant, and assuming uniform acceleration, how high a stack of hay is required for the crew member to survive the fall? - I'm completely lost. Where do I begin?

OpenStudy (shane_b):

This problem is interesting so I'll take a stab at it. Someone else should check it and see if it makes sense: First convert 127.6 km/hr to m/s. You should get 35.44 m/s. Next calculate the max acceleration the person can withstand (assuming downward is negative for this problem): \[(35)(-9.8m/s^2)=-343m/s^2\] So here's what we have so far:\[V_{final} = 0 m/s\]\[V_{initial} = -35.44 m/s\]\[a = -343 m/s^2\] Finally, here's the kinematic equation to use: \[V_f^2=V_i^2+2ad\] Rearrange it to solve for d: \[d=\frac{Vf^2-Vi^2}{2a}\]Plug in values to get the required height:\[d=\frac{(0m/s)^2-(-35.44m/s)^2}{2(-343m/s^2)}=1.831m\]

OpenStudy (shane_b):

Basically, if the height were any lower than 1.831m, you'd exceed 35g.

OpenStudy (anonymous):

Ah, now it makes sense! I was unsure as to whether I should use 127.6 km/hr as initial velocity, and it was weird trying to think about calculating a distance from the ground up instead of the opposite. I thought I needed a really complicated formula or something. Thank you so much!!